OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomore tickets totaled $400. How many sophomores bought tickets?

OpenStudy (karatechopper):

2.00f+3.00s=400

OpenStudy (karatechopper):

f+s=150

OpenStudy (karatechopper):

f=freshman s=sophmore

OpenStudy (karatechopper):

its a system of equations!

OpenStudy (karatechopper):

do you know how to solve these?

OpenStudy (anonymous):

no

OpenStudy (karatechopper):

ok then, well lets see what grade r u in?

OpenStudy (anonymous):

9

OpenStudy (karatechopper):

ok im in 7th grade:) now i know how well to explain, so a system of equations gives u 2 answers, always! u just have to solve for em

OpenStudy (karatechopper):

so looking at your problem, what do we have to solve for?

OpenStudy (anonymous):

2.00f+3.00(150 - f)=400 2 f + 450 -3f =400 -f = -50 f =50 and s = 100